# CommonPower: Supercharging Machine Learning for Smart Grids

### v1

@@ VII Conclusion
We introduce the Python library CommonPower – a tool for the design and benchmark creation of safe RL controllers for smart grids. It facilitates convenient modeling of power systems and maintains a symbolic system representation. The modular architecture, in combination with the symbolic representation, makes it possible to explicitly study different RL safeguarding strategies, while the built-in model predictive controller can readily serve as a credible baseline for performance evaluation. CommonPower enables the simulation of single-agent and multi-agent deployments via a unified interface as well as the combination of different controller types in the same system. The usage of possibly inaccurate forecasts is made explicit, enabling detailed studies on the impact of forecaster accuracy. Planned extensions of CommonPower are the implementation of further entity models, robust optimal control, hierarchical control, contingency constraints, further safeguarding approaches, and explicit modeling of energy trading.

### v3

@@ VII Conclusion
We introduce a comprehensive toolbox, CommonPower, for the study of safe RL controllers in smart grids. CommonPower offers convenient modeling of power systems and serves as a modular framework facilitating comparisons of different safeguarding approaches, levels of decentralization, controller types, and forecasters. Leveraging a symbolic representation of the system, CommonPower automatically extracts system constraints, which serves as the basis for, firstly, a robust model-based safeguarding framework that explicitly considers inaccurate forecasts and, secondly, the automatic derivation of a model predictive controller that can readily serve as a credible baseline for performance evaluation. Moreover, CommonPower enables the simulation of single-agent and multi-agent deployments via a unified interface as well as the combination of RL-based and non-RL controllers in the same system. Planned extensions of CommonPower are the implementation of further entity models, hierarchical control, contingency constraints, further safeguarding approaches, and explicit modeling of energy trading.